OpenStudy (anonymous):

you purchase a new car a $31,000 and the value decreases %30 each year. How many years will it take before your car is worth lass than $24,000? less than $5,000?

OpenStudy (anonymous):

i figured out the function of the value of the car would be 31,000(.70)^t

OpenStudy (anonymous):

a) 31,000(.70)^t = 24,000 b) 31,000(.70)^t = 5,000

OpenStudy (anonymous):

yea i got that but i dont know how to find t where its less than $5000

OpenStudy (anonymous):

take the log of both sides. log(31,000(.70)^t) = log(5,000) log(31,000)+log(.7^t) = log(5,000) t log(.7) = log(5,000)-log(31,000) t = (log(5,000)-log(31,000))/log(.7)

OpenStudy (anonymous):

ohh okay so u want to get the t on one side..so t will equal -.938 (rounded) but how will it be less than 5000 at negative.938 years?

OpenStudy (anonymous):

t should be 5.115

OpenStudy (anonymous):

t = (log(5,000)-log(31,000))/log(.7) = log(5/31)/log(.7)

OpenStudy (anonymous):

ohh okay thought it was 7 not .7 log srry k i get it know so that means at 5.115 years its value is less thank 5000?

OpenStudy (anonymous):

at 5.115 years its value is equal to 5000 beyond 5.115 years, its value will be less than 5000 so after 6 years, its value will be less than 5000

OpenStudy (anonymous):

okay thank u soo much...helps me 4 my upcoming exam

OpenStudy (anonymous):

:)